What is the distance between (6,-1)and (-3,7)?
Answer by checkley77(12569)

distance between (6,-1)and (-3,7)is:
for the x distance we have 6+3=9
for the y distance we have 7+1=8
thus we have a right triangle with 2 sides of 9 & 8 & we're looking for the hypotenuse using:
[a^2+b^2=c^2].
8^2+9^2=h^2
64+81=h^2
145=h^2
h=sqrt145
h=12.0416 is the distance between these 2 points.